# CECILIA: Gas-Phase Physical Conditions and Multi-Element Chemistry at Cosmic Noon

### v1

@@ 2.6 Combining MOSFIRE and NIRSpec Data
To make this comparison, we normalize the [O II] emission lines by the intensity of H or H observed in MOSFIRE and NIRSpec. This approach assumes that all line ratios (e.g., [O II]3727,29/H) remain constant between the NIRSpec and MOSFIRE observations of the same galaxy. However, such an assumption may be invalid if the different observations trace distinct components of the ionized gas. The archival MOSFIRE observations of the CECILIA galaxies are acquired at a variety of PA, with slits of width 07, and median extraction length 18 (A. L. Strom et al., 2017). The MOSFIRE slits are significantly larger than the MSA slits, which have width 020 and length ranging from 3-8 shutters (corresponding to 138-368) for the CECILIA galaxies. The MSA slits were selected based on the assigned PA range at time of observations (see discussion in A. L. Strom et al., 2023), and small differences in the positioning of the MSA slit relative to the MOSFIRE observations could produce different line ratios in the NIRSpec and MOSFIRE spectra.
@@
The stability of the CEL-to-H I line ratios is reassuring, although we note that a similar comparison cannot be made for the [O II] nebular lines. Normalizing the [O II] lines to H or H in MOSFIRE is sensitive to the relative flux calibration of different NIR observations and can be impacted by slit losses (see discussion and approach adopted for KBSS in A. L. Strom et al., 2017). While we acknowledge that slit losses and observational effects may alter the comparison of emission line ratios, the agreement observed in Figure 2 motivates us to use the relative emission line ratios in the NIRSpec and MOSFIRE data to carry out the following Te and abundance analysis.
@@ 2.7 Reddening Corrections
We correct for dust attenuation using the H and H recombination lines in the NIRSpec and MOSFIRE data. A measurement of the Balmer decrement in NIRSpec is possible in 7 galaxies, and the majority of the galaxies have H/H ratios available in the MOSFIRE data. We compare the observed flux ratio to the theoretical Case B H/H ratio at Te 1.25104 K and ne 300 cm-3, which is 2.82 using the atomic data of P. J. Storey & D. G. Hummer (1995). While the Balmer lines are sensitive to underlying stellar absorption, this is generally a small correction for H and H (6% to the H flux in KBSS, A. L. Strom et al., 2017), and the absorption component is accounted for in the SED continuum. The required to produce the observed Balmer decrement is computed assuming a N. A. Reddy et al. (2020) attenuation curve, which is calibrated on the emission line trends of galaxies at Cosmic Noon from the MOSFIRE Deep Evolution Field survey (MOSDEF, M. Kriek et al., 2015). The functional form of this attenuation curve is similar to the Galactic attenuation curve of J. A. Cardelli et al. (1989), although recent JWST observations suggest that the shape of the attenuation curve may evolve at z2 (V. Markov et al., 2024; R. L. Sanders et al., 2024b; N. A. Reddy et al., 2025).
While the Balmer decrement is unavailable in NIRSpec for most CECILIA targets, we can use the significant Paschen line detections in the deep G235M data along with H to measure . This is discussed further in Appendix B, although we note that these calculations rely on the shape of the N. A. Reddy et al. (2020) attenuation curve in an area where it is not well calibrated. As such, when reporting for the CECILIA galaxies we prioritize measured from the NIRSpec Balmer decrement, followed by the MOSFIRE Balmer decrement, and finally an average from the available Paschen lines in the G235M or G395M data. For galaxies RK120 and BX523, we use from the Paschen lines owing to uncertain MOSFIRE slit-loss corrections. The adopted in each galaxy is reported in Table 2. Emission line ratios most sensitive to include the auroral-to-nebular line ratios of [O II] and [S III] in addition to [O II]3727,29/H required to measure O+/H+. Systematic uncertainties in the attenuation curve for high-z galaxies could significantly bias the physical conditions and chemical abundances reported in the CECILIA galaxies. Therefore, we adopt an approach that uses neighboring H I lines for reddening-insensitive Te and abundance calculations, which is discussed in §3.2.1.
@@ 2.8 CECILIA Properties and Emission Line Trends
In Figure 3, we investigate whether the star-formation and ionization conditions of the CECILIA sample is representative of the parent sample, KBSS. In both panels, we denote the CECILIA galaxies as white circles, where the red circles represent the galaxies with at least one direct Te measurement. We plot the star-forming main sequence (SFMS), log(SFR) vs. log(M∗/M⊙), for KBSS (blue contours at the 5%, 25%, 50%, 75%, and 95% densities)444Note that the full KBSS distribution displays an artificial linear feature corresponding to an apparent maximum sSFR of . As noted in §2.2, this feature is due to the SED-fit SFRs being averaged over a yr timescale that is longer than the yr bursts allowed by our continuity SFH model. and CECILIA. Here, the SFR and M∗ for all galaxies are determined through BAGPIPES SED fitting to the available KBSS photometric data using non-parametric SFHs (see §2.2 and N. A. Korhonen Cuestas et al., 2025). The CECILIA galaxies are consistent with the SFMS obtained from the full KBSS sample, where the Te sample spans log(SFR) [0.22,1.69] and log(M∗/M⊙) [8.32,10.55], and is not significantly elevated in SFR relative to the parent sample.
@@
Plotted in the right panel of Figure 3 is the N2-BPT diagram (see J. A. Baldwin et al., 1981; S. Veilleux & D. E. Osterbrock, 1987), which uses the emission line ratios of [O III]/H and [N II]/H. The evolution of these line ratios in SFGs is sensitive to the metallicity of the ISM, the electron density, and the ionizing spectrum incident upon the gas (e.g., L. J. Kewley et al., 2013). All CECILIA galaxies plotted in this diagram have [N II]/H measured from NIRSpec, but NIRSpec [O III]/H ratios are only available in seven galaxies. For the remainder, we use the line ratios measured from the MOSFIRE spectra.
@@
The demarcation curves separating line ratios observed in local AGN (G. Kauffmann et al., 2003, dotted green line) and predicted for extreme starbursts (L. J. Kewley et al., 2001, blue dashed line) are plotted, as are the average line ratios measured in local SFGs from the Sloan Digital Sky Survey (SDSS, D. G. York et al., 2000, solid black) and KBSS sample by A. L. Strom et al. (2017, solid pink). The CECILIA line ratios agree with the KBSS line ratios in the N2-BPT diagram, indicating that CECILIA is a representative subsample of KBSS in terms of average ionization and metallicity. Furthermore, the line ratios are consistent with the general emission line trends at high-z, namely an offset to more intense CEL emission relative to z0 SFGs (see C. C. Steidel et al., 2014; A. E. Shapley et al., 2015; D. Kashino et al., 2017). We note that galaxies with low [O III]/H in KBSS were not prioritized for the CECILIA sample (see §2.1), as auroral line emission is prohibitively faint in metal rich galaxies with low-ionization conditions.
@@
In summary, the stellar and ionizing properties of the CECILIA galaxies are representative of the KBSS sample and they are in good agreement with the general trends observed in galaxies at Cosmic Noon. Crucially, we observe that direct Te measurements are possible in a diverse population of galaxies, including those on the SFMS at low and high stellar masses. The 20 galaxies with at least one Te measurement span the full range in log(M∗) and log(SFR) targeted in the CECILIA galaxies and have ionization conditions of log() [0.44, 0.87], and log() [0.76, 1.14]. Most prior Te measurements at high-z have been made in galaxies significantly above from the SFMS (D. Scholte et al., 2025), which is a consequence of the ISM conditions necessary for auroral line excitation. As Figure 3 demonstrates, it is possible to significantly detect faint auroral line emission in a variety of SFGs with deep NIRSpec observations, allowing us to characterize the physical conditions and metal production in the typical SFG at Cosmic Noon.
@@ 3 Physical Conditions and Chemical Abundances
| N+ | C. Froese Fischer & G. Tachiev (2004) | S. S. Tayal (2011) |
| O+ | C. Froese Fischer & G. Tachiev (2004) | R. Kisielius et al. (2009) |
| O2+ | C. Froese Fischer & G. Tachiev (2004) | P. J. Storey et al. (2014) |
| S+ | A. Irimia & C. Froese Fischer (2005) | S. S. Tayal & O. Zatsarinny (2010) |
| S2+ | C. Froese Fischer et al. (2006) | C. E. Hudson et al. (2012) |
| Ar2+ | C. Mendoza & C. J. Zeippen (1983) | J. M. Munoz Burgos et al. (2009) |
With the combination of the NIRSpec and MOSFIRE data, it is possible to determine the gas-phase physical conditions and chemical composition of the ISM in the CECILIA galaxies. We employ the PyNeb package (V. Luridiana et al., 2015) for these calculations, and we summarize the assumed atomic data in Table 1. We consider the ISM to be composed of three zones, separated based on the ionization potentials of the metals within the gas. A two-zone model (e.g., C. Esteban et al., 2009) is often invoked when measuring the ionic abundances of O, where the conditions of the [O II] and [O III] emitting gas define the low- and high-ionization zones, respectively. These zones contain other metal ions with similar ionization potentials: N+, O+, and S+ reside in the low-ionization zone, while the high-ionization zone contains O2+. However, S2+ and Ar2+ have ionization potentials that do not align with either O+ or O2+, which has motivated the use of a third, intermediate-ionization zone (e.g., D. R. Garnett, 1992). Indeed, spectroscopic observations of local star-forming nebulae indicate that the Te trends in the intermediate-ionization zone are distinct from the Te in the simple two-zone model (R. C. Kennicutt et al., 2003; K. V. Croxall et al., 2016; N. S. J. Rogers et al., 2021; J. E. Méndez-Delgado et al., 2023), requiring either a directly measured or inferred Te to accurately measure the abundances of S2+ and Ar2+. Given our numerous detections of the [S III] auroral line, we can constrain the Te[S III] and ionic abundances of S2+ and Ar2+ in this intermediate-ionization zone within the CECILIA galaxies.
@@ 3.1 Electron Densities
There are two ne-sensitive doublets routinely detected in the CECILIA spectroscopic data. ne[S II] is measured from [S II]6718,33 in the NIRSpec data, and ne[O II] is constrained using [O II]3727,29 from the higher resolution MOSFIRE rest-optical spectra (R 3300 in band). This allows us to compare the ne inferred from different low-ionization species, which would not be possible from the NIRSpec medium-resolution gratings (R 1000) as the [O II] strong lines require R 1350 to resolve. This is true even for compact targets where the effective resolution of the medium gratings is 1000 (A. de Graaff et al., 2024a), requiring the use of high-resolution gratings to reliably separate the [O II] lines and constrain ne[O II] (e.g., S. Li et al., 2025; B. Welch et al., 2025). Owing to the observational configuration of KBSS, the [O II] strong lines are generally observed in band, which has many bright OH lines. Since ne measurements require an accurate fit to the shape of the emission line doublet, we use the FWHM of [O III]5008 to constrain the FWHM of the [O II] lines (see §2.5) and inspect all [O II] emission lines in the MOSFIRE spectra to ensure that both lines are reliably fit and uncontaminated from sky line residuals. In instances where the doublet cannot be reliably fit (five galaxies total), we do not attempt a ne[O II] measurement.
@@
Other optical density diagnostics from scarce ions such as Cl2+ and Ar3+ are either not covered in the CECILIA observations or are too faint to detect even in the ultra-deep G235M observations. The density-sensitive [Cl III] lines have yet to be detected in JWST spectroscopy of any high-z SFG; while the [Ar IV] doublet has been reported in a handful of high-z nebulae (N. S. J. Rogers et al., 2024; T. Morishita et al., 2025; B. Welch et al., 2025), the neighboring He I 4715 emission line must be properly modeled to accurately measure ne[Ar IV]. The CECILIA observations do not cover the near-UV C III 1907,09 doublet, which has provided a tracer of the electron density in the high-ionization ISM of high-z SFGs (e.g., M. W. Topping et al., 2024, 2025a). While this line doublet is generally too faint to detect in the Keck/LRIS observations of individual KBSS galaxies, a ne[C III] of 1470660 cm-3 was reported in a stack of 30 KBSS galaxies (C. C. Steidel et al., 2016).
@@
For ne calculations, we use the PyNeb getTemDen function, which takes the ne-sensitive line intensity ratio and an assumed gas temperature and returns the density required to produce the observed line ratio. While the intensity of an individual CEL in the doublet is sensitive to both Te and ne, the line ratio mitigates much of the Te dependence because the CELs have similar excitation energies. To calculate ne, we assume a Te of 1.25104 K based on the initial temperature measurements in the CECILIA galaxies (A. L. Strom et al., 2023; N. S. J. Rogers et al., 2024). To estimate uncertainties on ne, we resample the line intensity ratio 1500 times from a normal distribution with standard deviation equal to the uncertainty on the line ratio, then recalculate ne for the distribution excluding any ratios outside the range permitted by the atomic data. We calculate the asymmetric errors on ne as the 16th and 84th percentiles of the distribution.
@@
The lower panel of Figure 4 plots the 15 simultaneous ne[O II] and ne[S II] measured from the MOSFIRE and NIRSpec observations, respectively, as red circles. The average densities measured in the CECILIA galaxies with both ne[S II] and ne[O II] (plotted as a dark red star) are: ne[S II]26744 cm-3 and ne[O II]39573 cm-3, where the uncertainties are estimated via bootstrap resampling the average statistic. The average [O II] density is slightly elevated compared to the [S II] density, although the individual ne measurements scatter around the one-to-one line (dotted line). This trend may be expected for two ions originating in the same volume of the ISM, although the ionization potential of S+ is such that this ion may also reside outside the H II region in the photodissociation region (PDR) and diffuse ionized gas (DIG).
@@
The typical electron density of the DIG is on the order of 0.01-0.1 cm-3 (R. J. Reynolds, 1991; E. M. Berkhuijsen & A. Fletcher, 2008), which is much lower than the ne within the H II regions. A substantial DIG contribution to the [S II] intensity could bias ne[S II] low relative to ne[O II], which is measured within the volume of the H II region. However, M. S. Oey et al. (2007) found a decreasing DIG fraction in local galaxies with high H surface brightness, such that galaxies with high SFR surface densities are expected to have minimal DIG contributions to the integrated galaxy spectrum (see also K. Zhang et al., 2017). A. E. Shapley et al. (2019) used the SFR surface densities in the MOSDEF sample to predict the DIG fraction as a function of z (their Equation 2, based on fit by R. L. Sanders et al., 2017), finding that the DIG contribution is negligible for galaxies at z2.3.
@@
The slightly elevated ne[O II] compared to ne[S II] in the CECILIA galaxies, therefore, may be related to a marginal contribution of the DIG to the [S II] line intensities, where the density measured from these CELs is more reflective of the ne in the H II regions. Alternatively, the DIG at z2 may exhibit elevated electron density relative to the DIG in local SFGs. If the DIG and H II regions are at comparable densities, then ne[S II] is unbiased by the inclusion of the DIG in the integrated spectrum. The agreement between ne[S II] and ne[O II] could also be related to the size and position of the MSA slits: if the small slits target the brightest H II regions and do not incorporate a significant DIG component in the integrated spectrum, then ne[S II] should better reflect the H II region density, similar to ne[O II]. In summary, if low-density DIG is present in the integrated spectra CECILIA galaxies, then its contribution to the [S II] line doublet only slightly decreases the average ne[S II] relative to ne[O II], or the DIG is described by marginally higher ne that does not produce a substantial bias to ne[S II].
@@
Overall, the density trends measured in the CECILIA galaxies are in good agreement with prior estimates of ISM ne at Cosmic Noon. Large ground-based spectroscopic surveys of galaxies at Cosmic Noon have consistently found that the average ne[O II] and ne[S II] range from 150-300 cm-3 (R. L. Sanders et al., 2016; A. L. Strom et al., 2017; R. L. Davies et al., 2021) These findings have been corroborated with JWST observations (e.g., Y. Isobe et al., 2023; S. Li et al., 2025; T. M. Stanton et al., 2025). Recently, M. W. Topping et al. (2025a) reported a gradual increase in ne[S II] with z from 51 galaxies at z1.38 observed as part of the AURORA survey (A. E. Shapley et al., 2024). The average ne[S II]268 cm-3 measured from the 37 AURORA galaxies at z2 is in excellent agreement with the average reported in Figure 4.
@@
The density trends observed at Cosmic Noon are relatively uncommon in local SFGs, where ne[S II] is typically 100 cm-3 (M. Kaasinen et al., 2017). As a comparison, we plot the ne[O II] and ne[S II] trends in a sample of 55 metal-poor, Extreme Emission Line Galaxies (EELGs) at z0 (Rogers et al., in preparation) as blue squares in Figure 4. The average ne[O II] and ne[S II] in this sample are 18021 cm-3 and 13618 cm-3, respectively, indicating that these two densities are closely related (see also J. E. Méndez-Delgado et al., 2023). The median z0 densities are in good agreement with the ne[O II] and ne[S II] observed in samples of other z0 EELGs (P. Senchyna et al., 2017; M. Mingozzi et al., 2022), some of which are selected to have high SFRs typical of those in SFGs at Cosmic Noon. However, the average densities measured in the z0 EELGs are approximately a factor of two less than the average ISM density observed in the CECILIA galaxies. The offset is observed when considering all 23 ne[S II] measurements in the CECILIA sample (i.e., including eight galaxies without ne[O II]), where the average is 21234 cm-3. This is highlighted in the top panel of Figure 4, which plots the histogram distribution of log(ne[S II]) for the full sample. A portion of the CECILIA galaxies are described by ne[S II]100 cm-3 similar to the local EELGs, but the peak of the distribution is offset to higher ne[S II].

### v2

@@ 2.6 Combining MOSFIRE and NIRSpec Data
To make this comparison, we normalize the [O II] emission lines by the intensity of H or H observed in MOSFIRE and NIRSpec. This approach assumes that all line ratios (e.g., [O II]3727,29/H) remain constant between the NIRSpec and MOSFIRE observations of the same galaxy. However, such an assumption may be invalid if the different observations trace distinct components of the ionized gas. The archival MOSFIRE observations of the CECILIA galaxies are acquired at a variety of PA, with slits of width 07, and median extraction length 18 (strom2017). The MOSFIRE slits are significantly larger than the MSA slits, which have width 020 and length ranging from 3-8 shutters (corresponding to 138-368) for the CECILIA galaxies. The MSA slits were selected based on the assigned PA range at time of observations (see discussion in strom2023), and small differences in the positioning of the MSA slit relative to the MOSFIRE observations could produce different line ratios in the NIRSpec and MOSFIRE spectra.
@@
The stability of the CEL-to-H I line ratios is reassuring, although we note that a similar comparison cannot be made for the [O II] nebular lines. Normalizing the [O II] lines to H or H in MOSFIRE is sensitive to the relative flux calibration of different NIR observations and can be impacted by slit losses (see discussion and approach adopted for KBSS in strom2017). While we acknowledge that slit losses and observational effects may alter the comparison of emission line ratios, the agreement observed in Figure 2 motivates us to use the relative emission line ratios in the NIRSpec and MOSFIRE data to carry out the following Te and abundance analysis.
@@ 2.7 Reddening Corrections
We correct for dust attenuation using the H and H recombination lines in the NIRSpec and MOSFIRE data. For the galaxies studied in this work, a measurement of the Balmer decrement at S/N 5, the recommended S/N for measuring in strom2017, is available in the NIRSpec and MOSFIRE data for 7 and 16 galaxies, respectively. We compare the observed flux ratio to the theoretical Case B H/H ratio at Te 1.25104 K and ne 300 cm-3, which is 2.82 using the atomic data of storey1995. While the Balmer lines are sensitive to underlying stellar absorption, this is generally a small correction for H and H (6% to the H flux in KBSS, strom2017), and the absorption component is accounted for in the SED continuum. The required to produce the observed Balmer decrement is computed assuming a reddy2020 attenuation curve, which is calibrated on the emission line trends of galaxies at Cosmic Noon from the MOSFIRE Deep Evolution Field survey (MOSDEF, kriek2015). The functional form of this attenuation curve is similar to the Galactic attenuation curve of cardelli1989, although recent JWST observations suggest that the shape of the attenuation curve may evolve at z2 (markov2024; sanders2024; reddy2025). When H is available in the G235M data for galaxies at z2.82, we derive as a weighted average of the reddening determined from the H/H and H/H ratios.
While the Balmer decrement is unavailable in NIRSpec for most CECILIA targets, we can use the significant Paschen line detections in the deep G235M data along with H to measure . This is discussed further in Appendix B, although we note that these calculations rely on the shape of the reddy2020 attenuation curve in an area where it is not well calibrated. As such, when reporting for the CECILIA galaxies we prioritize measured from the NIRSpec Balmer decrement, followed by the MOSFIRE Balmer decrement, and finally an average from the available Paschen lines in the G235M or G395M data relative to H. For six galaxies with measurements from the Balmer decrement and multiple Paschen lines, we find that from the Balmer lines is, on average, larger by 0.09 mag. This is opposite to the average trends measured in galaxies at z1.5 by reddy2025, and we note that a lack of high-order Balmer lines that could better constrain in the rest-optical may contribute to this offset in the CECILIA galaxies. For galaxies RK120 and BX523, we use from the Paschen lines owing to uncertain MOSFIRE slit-loss corrections.
The adopted in each galaxy is reported in Table 2. Emission line ratios most sensitive to include the auroral-to-nebular line ratios of [O II] and [S III] in addition to [O II]3727,29/H required to measure O+/H+. Systematic uncertainties in the attenuation curve for high-z galaxies could significantly bias the physical conditions and chemical abundances reported in the CECILIA galaxies. Therefore, we adopt an approach that uses neighboring H I lines for reddening-insensitive Te and abundance calculations, which is discussed in §3.2.1.
@@ 2.8 CECILIA Properties and Emission Line Trends
In Figure 3, we investigate whether the star-formation and ionization conditions of the CECILIA sample is representative of the parent sample, KBSS. In both panels, we denote the CECILIA galaxies as white circles, where the red circles represent the galaxies with at least one direct Te measurement. We plot the star-forming main sequence (SFMS), log(SFR) vs. log(M∗/M⊙), for KBSS (blue contours at the 5%, 25%, 50%, 75%, and 95% densities)444Note that the full KBSS distribution displays an artificial linear feature corresponding to an apparent maximum sSFR of . As noted in §2.2, this feature is due to the SED-fit SFRs being averaged over a yr timescale that is longer than the yr bursts allowed by our continuity SFH model. and CECILIA. Here, the SFR and M∗ for all galaxies are determined through BAGPIPES SED fitting to the available KBSS photometric data using non-parametric SFHs (see §2.2 and korhonen-cuestas2025). The CECILIA galaxies are consistent with the SFMS obtained from the full KBSS sample, where the Te sample spans log(SFR) [0.22,1.69] and log(M∗/M⊙) [8.32,10.55], and is not significantly elevated in SFR relative to the parent sample.
We also provide fits to the SFMS at Cosmic Noon from speagle2014, popesso2023, and clarke2024. We have selected the SFMS most appropriate for comparison to the CECILIA sample: for speagle2014 and popesso2023, the plotted relations are determined at z2.5, while the clarke2024 relation is derived from SFGs at 1.4z2.7 with SED-based SFRs. In general, the CECILIA galaxies with Te measurements scatter between the various measurements of the SFMS at Cosmic Noon. We note that using the H luminosities measured from the G235M spectra produce SFRs that are, on average, 0.26 dex larger than measured from the SED fits. This reflects recent ( 10 Myr) star formation within the CECILIA galaxies that the SED fitting is insensitive to on account of the line-corrected photometry.
@@
Plotted in the right panel of Figure 3 is the N2-BPT diagram (see baldwin1981; veilleux1987), which uses the emission line ratios of [O III]/H and [N II]/H. The evolution of these line ratios in SFGs is sensitive to the metallicity of the ISM, the electron density, and the ionizing spectrum incident upon the gas (e.g., kewley2013). All CECILIA galaxies plotted in this diagram have [N II]/H measured from NIRSpec, but NIRSpec [O III]/H ratios are only available in seven galaxies. For the remainder, we use the line ratios measured from the MOSFIRE spectra.
@@
The demarcation curves separating line ratios observed in local AGN (kauffmann2003, dotted green line) and predicted for extreme starbursts (kewley2001, blue dashed line) are plotted, as are the average line ratios measured in local SFGs from the Sloan Digital Sky Survey (SDSS, york2000, solid black) and KBSS sample by strom2017. The CECILIA line ratios agree with the KBSS line ratios in the N2-BPT diagram, indicating that CECILIA is a representative subsample of KBSS in terms of average ionization and metallicity. Furthermore, the line ratios are consistent with the general emission line trends at high-z, namely an offset to more intense CEL emission relative to z0 SFGs (see steidel2014; shapley2015; kashino2017). We note that galaxies with low [O III]/H in KBSS were not prioritized for the CECILIA sample (see §2.1), as auroral line emission is prohibitively faint in metal rich galaxies with low-ionization conditions.
@@
In summary, the stellar and ionizing properties of the CECILIA galaxies are representative of the KBSS sample and they are in good agreement with the general trends observed in galaxies at Cosmic Noon. Crucially, we observe that direct Te measurements are possible in a diverse population of galaxies, including those on the SFMS at low and high stellar masses. The 20 galaxies with at least one Te measurement span the full range in log(M∗) and log(SFR) targeted in the CECILIA galaxies and have ionization conditions of log() [0.44, 0.87], and log() [0.76, 1.14]. Most prior Te measurements at high-z have been made in galaxies significantly above from the SFMS (scholte2025), which is a consequence of the ISM conditions necessary for auroral line excitation. As Figure 3 demonstrates, it is possible to significantly detect faint auroral line emission in a variety of SFGs with deep NIRSpec observations, allowing us to characterize the physical conditions and metal production in the typical SFG at Cosmic Noon.
@@ 3 Physical Conditions and Chemical Abundances
| N+ | froesefischer2004 | tayal2011 |
| O+ | froesefischer2004 | kisielius2009 |
| O2+ | froesefischer2004 | storey2014 |
| S+ | irimia2005 | tayal2010 |
| S2+ | froesefischer2006 | hudson2012 |
| Ar2+ | mendoza1983 | munozburgos2009 |
With the combination of the NIRSpec and MOSFIRE data, we can now determine the gas-phase physical conditions and chemical composition of the ISM in the CECILIA galaxies. We employ PyNeb v1.1.18 (luridiana2015; morisset2020) for these calculations, and we summarize the assumed atomic data in Table 1. We consider the ISM to be composed of three zones, separated based on the IPs of the metals within the gas. A two-zone model (e.g., esteban2009) is often invoked when measuring the ionic abundances of O, where the conditions of the [O II] and [O III] emitting gas define the low- and high-ionization zones, respectively. These zones contain other metal ions with similar IPs: N+, O+, and S+ reside in the low-ionization zone (10.1 eV IP 23.4 eV), while the high-ionization zone contains O2+ (IP 35.1 eV). However, S2+ and Ar2+ have IPs that do not align with either O+ or O2+, which has motivated the use of a third, intermediate-ionization zone (e.g., garnett1992) for ions with 23.4 eV IP 35.1 eV. Indeed, spectroscopic observations of local star-forming nebulae indicate that the Te trends in the intermediate-ionization zone are distinct from the Te in the simple two-zone model (kennicutt2003; croxall2016; rogers2021; mendez-delgado2023), requiring either a directly measured or inferred Te to accurately measure the abundances of S2+ and Ar2+. Given our numerous detections of the [S III] auroral line, we can constrain the Te[S III] and ionic abundances of S2+ and Ar2+ in this intermediate-ionization zone within the CECILIA galaxies.
@@ 3.1 Electron Densities
There are two ne-sensitive doublets routinely detected in the CECILIA spectroscopic data. ne[S II] is measured from [S II]6718,33 in the NIRSpec data, and ne[O II] is constrained using [O II]3727,29 from the higher resolution MOSFIRE rest-optical spectra (R 3300 in band). This allows us to compare the ne inferred from different low-ionization species, which would not be possible from the NIRSpec medium-resolution gratings (R 1000) as the [O II] strong lines require R 1350 to resolve. This is true even for compact targets where the effective resolution of the medium gratings is 1000 (degraaf2024b), requiring the use of high-resolution gratings to reliably separate the [O II] lines and constrain ne[O II] (e.g., li2025; welch2025). Owing to the observational configuration of KBSS, the [O II] strong lines are generally observed in band, which has many bright OH lines. Since ne measurements require an accurate fit to the shape of the emission line doublet, we use the FWHM of [O III]5008 to constrain the FWHM of the [O II] lines (see §2.5) and inspect all [O II] emission lines in the MOSFIRE spectra to ensure that both lines are reliably fit and uncontaminated from sky line residuals. In instances where the doublet cannot be reliably fit (five galaxies total), we do not attempt a ne[O II] measurement.
@@
Other optical density diagnostics from scarce ions such as Cl2+ and Ar3+ are either not covered in the CECILIA observations or are too faint to detect even in the ultra-deep G235M observations. The density-sensitive [Cl III] lines have yet to be detected in JWST spectroscopy of any high-z SFG; while the [Ar IV] doublet has been reported in a handful of high-z nebulae (rogers2024; morishita2025; welch2025), the neighboring He I 4715 emission line must be properly modeled to accurately measure ne[Ar IV]. The CECILIA observations do not cover the near-UV [C III]1907,09 doublet, which has provided a tracer of the electron density in the high-ionization ISM of high-z SFGs (e.g., topping2024; topping2025). While this line doublet is generally too faint to detect in the Keck/LRIS observations of individual KBSS galaxies, a ne[C III] of 1470660 cm-3 was reported in a stack of 30 KBSS galaxies (steidel2016).
@@
For ne calculations, we use the PyNeb getTemDen function, which takes the ne-sensitive line intensity ratio and an assumed gas temperature and returns the density required to produce the observed line ratio. While the intensity of an individual CEL in the doublet is sensitive to both Te and ne, the line ratio mitigates much of the Te dependence because the CELs have similar excitation energies. To calculate ne, we assume a Te of 1.25104 K based on the initial temperature measurements in the CECILIA galaxies (strom2023; rogers2024). To estimate uncertainties on ne, we resample the line intensity ratio 1500 times from a normal distribution with standard deviation equal to the uncertainty on the line ratio, then recalculate ne for the distribution excluding any ratios outside the range permitted by the atomic data. We calculate the asymmetric errors on ne as the 16th and 84th percentiles of the distribution.
@@
The lower panel of Figure 4 plots the 15 simultaneous ne[O II] and ne[S II] measured from the MOSFIRE and NIRSpec observations, respectively, as red circles. The average densities measured in the CECILIA galaxies with both ne[S II] and ne[O II] (plotted as a dark red star) are: ne[S II]26744 cm-3 and ne[O II]38670 cm-3, where the uncertainties are estimated via bootstrap resampling the average statistic. The average [O II] density is slightly elevated compared to the [S II] density, although the individual ne measurements scatter around the one-to-one line (dotted line). This trend may be expected for two ions originating in the same volume of the ISM, although the IP of S+ is such that this ion may also reside outside the H II region in the photodissociation region (PDR) and diffuse ionized gas (DIG).
@@
The typical electron density of the DIG is on the order of 0.01-0.1 cm-3 (reynolds1991; berkhuijsen2008), which is much lower than the ne within the H II regions. A substantial DIG contribution to the [S II] intensity could bias ne[S II] low relative to ne[O II], which is measured within the volume of the H II region. However, oey2007 found a decreasing DIG fraction in local galaxies with high H surface brightness, such that galaxies with high SFR surface densities are expected to have minimal DIG contributions to the integrated galaxy spectrum (see also zhang2017). shapley2019 used the SFR surface densities in the MOSDEF sample to predict the DIG fraction as a function of z (their Equation 2, based on fit by sanders2017), finding that the DIG contribution is negligible for galaxies at z2.3.
@@
The slightly elevated ne[O II] compared to ne[S II] in the CECILIA galaxies, therefore, may be related to a marginal contribution of the DIG to the [S II] line intensities, where the density measured from these CELs is more reflective of the ne in the H II regions. Alternatively, the DIG at z2 may exhibit elevated electron density relative to the DIG in local SFGs. If the DIG and H II regions are at comparable densities, then ne[S II] is unbiased by the inclusion of the DIG in the integrated spectrum. The agreement between ne[S II] and ne[O II] could also be related to the size and position of the MSA slits: if the small slits target the brightest H II regions and do not incorporate a significant DIG component in the integrated spectrum, then ne[S II] should better reflect the H II region density, similar to ne[O II]. In summary, if low-density DIG is present in the integrated spectra of the CECILIA galaxies, then its contribution to the [S II] line doublet only slightly decreases the average ne[S II] relative to ne[O II], or the DIG is described by marginally higher ne that does not produce a substantial bias to ne[S II].
@@
Overall, the density trends measured in the CECILIA galaxies are in good agreement with prior estimates of ISM ne at Cosmic Noon. Large ground-based spectroscopic surveys of galaxies at Cosmic Noon have consistently found that the average ne[O II] and ne[S II] range from 150-300 cm-3 (sanders2016; strom2017; davies2021) These findings have been corroborated with JWST observations (e.g., isobe2023; li2025; stanton2025). Recently, topping2025 reported a gradual increase in ne[S II] with z from 51 galaxies at z1.38 observed as part of the AURORA survey (shapley2024). The average ne[S II]268 cm-3 measured from the 37 AURORA galaxies at z2 is in excellent agreement with the average reported in Figure 4.
@@
The density trends observed at Cosmic Noon are relatively uncommon in local SFGs, where ne[S II] is typically 100 cm-3 (kaasinen2017). As a comparison, we plot the ne[O II] and ne[S II] trends in a sample of 55 metal-poor, Extreme Emission Line Galaxies (EELGs) at z0 (Rogers et al., in preparation) as blue squares in Figure 4. The average ne[O II] and ne[S II] in this sample are 18021 cm-3 and 13618 cm-3, respectively, indicating that these two densities are closely related (see also mendez-delgado2023). The median z0 densities are in good agreement with the ne[O II] and ne[S II] observed in samples of other z0 EELGs (senchyna2017; mingozzi2022), some of which are selected to have high SFRs typical of those in SFGs at Cosmic Noon. However, the average densities measured in the z0 EELGs are approximately a factor of two less than the average ISM density observed in the CECILIA galaxies. The offset is observed when considering all 23 ne[S II] measurements in the CECILIA sample (i.e., including eight galaxies without ne[O II]), where the average is 21234 cm-3. This is highlighted in the top panel of Figure 4, which plots the histogram distribution of log(ne[S II]) for the full sample. A portion of the CECILIA galaxies are described by ne[S II]100 cm-3 similar to the local EELGs, but the peak of the distribution is offset to higher ne[S II].